OpenStudy (anonymous):

What is the value of the series: sum of (3^n+1)/(4^n) for n=1 to infinity.I will also write the equation w/ the eqution box

OpenStudy (anonymous):

teh equation box doesnt work

OpenStudy (anonymous):

it keeps saying math procesing error

OpenStudy (anonymous):

\[sum_{n=1}^{infty} (3^n+1)/ (4^n)\]

OpenStudy (anonymous):

\[\sum_{n=1}^{\infty}\frac{3^n+1}{4^n}\] like that?

OpenStudy (anonymous):

oh well even if you refresh?

OpenStudy (anonymous):

let me try

OpenStudy (anonymous):

oh i see it now

OpenStudy (anonymous):

no teh series is from n=1 to infinity numerator - 3^ (n+1) .....the n+1 are together denominator - 4^n

OpenStudy (anonymous):

or maybe it is \[\sum_{n=1}^{\infty}\frac{3^{n+1}}{4^n}\]

OpenStudy (anonymous):

yes that correct

OpenStudy (anonymous):

ok lets pull out a 3

OpenStudy (anonymous):

\[3\sum_{n=1}^{\infty}\frac{3^{n}}{4^n}\]

OpenStudy (anonymous):

what happedn to teh 1 that is with n+1

OpenStudy (anonymous):

\[3\sum_{k=1}^{\infty}(\frac{3}{4})^n\]

OpenStudy (anonymous):

i factored out a 3, precisely so the exponents would match

OpenStudy (anonymous):

oh because 3^n * 3 is 3^n+1 right

OpenStudy (anonymous):

and now it is easy enough right?

OpenStudy (anonymous):

so now u use teh geometric series test right, with \[\left| 3/4 \right| < 1 \] so it converges

OpenStudy (anonymous):

sure does

OpenStudy (anonymous):

are you summing from 1 to infinity or from zero to infinity?

OpenStudy (anonymous):

from 1 to infinity

OpenStudy (anonymous):

ohh so if its from one teh exponent has to be n-1 though!

OpenStudy (anonymous):

then the answer is \[\frac{\frac{3}{4}}{1-\frac{3}{4}}=\frac{\frac{3}{4}}{\frac{1}{4}}=3\]

OpenStudy (anonymous):

oh and don't forget to multiply by the 3 out front, so your real answer is 9

OpenStudy (anonymous):

wait why is a = 3/4

OpenStudy (anonymous):

you are startinng at n = 1, not zero

OpenStudy (anonymous):

so when u start from n=1, the a value is equal to the r value?

OpenStudy (anonymous):

so you have \[\frac{3}{4}+(\frac{3}{4})^2+(\frac{3}{4})^3+...\] \[\frac{3}{4}(1+\frac{3}{4}+(\frac{3}{4})^2+..\]

OpenStudy (anonymous):

so it looks like \[a+ar+ar^2+ar^3+...\] where \[a=\frac{3}{4}, r=\frac{3}{4}\] and you can use \[\frac{a}{1-r}\]

OpenStudy (anonymous):

so will it be safe to assume that when it is from n=1and its in the form ar^n, then the a and r value are the same

OpenStudy (anonymous):

i guess. lets look at is a simpler way if it was \[\sum_{n=0}^{\infty}(\frac{3}{4})^n\] you would get \[\frac{1}{1-\frac{3}{4}}=4\] but since you are starting at 1, you can subtract 1 from the result and get 3

OpenStudy (anonymous):

omg that makes perfect sense now!! =D

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

i did not even think of that

OpenStudy (anonymous):

don't forget to multiply by 3 to get your real answer good luck

OpenStudy (anonymous):

i cant believe seris are soo easy yet sometimes i dont see teh things right before my eyes and THANKS ALOT!!

OpenStudy (anonymous):

yw (alot)

OpenStudy (anonymous):

LOL

OpenStudy (anonymous):

wait can u please help me with one more problem when i post it